# bibliography.bib

### v1

@@ 4 Loss functions and the associated training methods
Having seen the expressivity of the ReLU NN and of the FKS and the formal equivalence of the (shallow) ReLU NN and the FKS approximation in Section 2, we consider different loss functions in this section which can be used to train these to find the best approximation of the given function $u$. We will describe several loss functions approximating the $L^{2}_{2}$ error, which, whilst formally equivalent, lead to different training behaviours. The training will be done through a (first order) gradient-based approach and using the Adam optimiser with a suitable initial start. This procedure is chosen as it is a very common way of training a machine learning network.
4.1 General loss functions for the ReLU NN, the linear spline approximations, and their training
To unify notation, we denote the ReLU NN or the FKS approximation by $y(x)\colon[0,1]\to\mathbb{R}$ and the associated parameter values by $\bm{\Theta}$. Further, we denote the target function by $u\colon[0,1]\to\mathbb{R}$. We consider the loss function $\mathcal{L}_{2}^{2}$ based directly on the mean-squared approximation error of the NN. As both $y(x)$ and $u(x)$ are given for all $x\in[0,1]$ we have that the $\mathcal{L}^{2}_{2}$-error is given by
@@ 4.1 General loss functions for the ReLU NN, the linear spline approximations, and their training
In practice, we cannot evaluate (14) for a general function and we have to consider various approximations to it. For this, we consider quadrature points $\{x_{k}\}_{k=0}^{s-1}\subset[0,1]$ for some large parameter $\0s\in\mathbb{N}\backslash\{0\}$ which can either be randomly generated with uniform distribution or regularly distributed over the domain $[0,1]$.
Assume that points $\{x_{k}\}_{k=0}^{s-1}\subset[0,1]$ are ordered so that $x_{i}<x_{i+1}$. An approximation to $\mathcal{L}_{2}^{2}$ in (14) is given by
$L_{2}^{2}(\bm{\Theta})=\sum_{i=0}^{s-1}(x_{i+1}-x_{i})\Big{(}y(x_{i},\bm{\Theta})-u(x_{i})\Big{)}^{2}.$ (15)
@@
Note that minimising (15) is a highly non-convex problem, and the minimisation process generally leads to sub-optimal solutions. A similar loss function is often used in the Deep Ritz Method (DRM) [E_DeepRitz]. Similarly to the DRM, we will consider both the case where $\{x_{k}\}_{k=0}^{s-1}$ are sampled at each iteration or fixed during the training.
The purpose of the training of the ReLU NN or the FKS is to vary the parameters in $\bm{\Theta}$ systematically to reduce the loss function. Typically this is done using the Adam optimiser. The parameters in $\bm{\Theta}$ are updated. Note that the iterates of $\bm{\Theta}$ depend on the choice of the loss function, the optimisation method used and the initial values of $\bm{\Theta}$. Note that the minimisation of $L_{2}^{2}$ in (15) reduces to the pointwise error at sample points $\{x_{i}\}_{i=0}^{s-1}$ and neglects global properties of the function, such as its curvature. This is important when considering equidistributed knots in regions of highest curvature, which contribute to the interpolation error significantly.
4.2 Loss functions for the interpolating FKS
@@
Whilst the interpolating FKS $\Pi_{1}u$ is sub-optimal as a general approximating function, it can be considered as an excellent first guess for general optimisation procedures. Finding the interpolating FKS only involves determining the knots $k_{i}$ for $i=1,\ldots,N-1$ which leads to a simplified approximation process approach for choosing the optimal knots which can then be combined with the notion of equidistribution. Having found the knots, the problem of finding the weights of the optimal FKS is a well conditioned linear problem.
4.2.1 The equidistribution loss function
@@
$y(x,\bm{\Theta})-u(x)\approx\frac{1}{2}(x-k_{i})(k_{i+1}-x)u^{\prime\prime}(x_{i+1/2})$ (16)
for any $x\in[k_{i},k_{i+1}]$ and some $x_{i+1/2}\in(k_{i},k_{i+1})$. Note that this result even holds for the function $u(x)=x^{2/3}$ for which $u^{\prime\prime}(0)$ is singular. This yields
@@
$L_{I}(\bm{\Theta})=\frac{1}{120}\sum_{i=0}^{N-2}\;(k_{i+1}-k_{i})^{5}\;|u^{\prime\prime}(k_{i+1/2})|^{2}$ (18)
where $k_{i+1/2}=k_{i}+k_{i+1}$ and we obtain $\mathcal{L}_{2}^{2}(\bm{\Theta})\approx L_{I}(\bm{\Theta})$. Note that $L_{I}$ only depends on the free knot locations for the interpolating FKS, and thus can be used to train the knots. However, unlike other loss functions $L_{I}$ requires knowledge of $u^{\prime\prime}$ rather than point values of $u$ and we will assume for the moment that $u^{\prime\prime}$ is known.
A powerful result in the form of the equidistribution principle, first introduced by de Boor [deBoor1973] for solving boundary value problems for ordinary differential equations, gives a way of finding $\bm{\Theta}$ so that $L_{I}$ in (18) is minimised. More precisely, minimising $L_{I}$ in (18) can be expressed by equidistributing the error over each cell which results in the following lemma (see [HuangRussel, Chapter 2]):
Lemma 1 (Equidistribution)
The loss $L_{I}$ in (18) is minimised over all choices of the knots $k_{i}$ if
@@
The equidistribution principle in Lemma 1 replaces a non-convex optimisation problem by one with better convexity properties. A set of knot points satisfying these conditions is said to be equidistributed. The degree of equidistribution can then be used as a quality measure of the solution.
The algebraic equations (20) can then be solved directly [deBoor1974], or iteratively using moving mesh techniques such as [Huang1994] for example. As an alternative, we propose training the interpolating FKS directly using an optimisation approach (such as Adam), but enforcing the equidistribution condition (20) directly through an equidistribution-based loss function.
To do this, we compute $\rho_{i+1/2}$ from knots $k_{i}$ by (19) and set $\sigma$ as their mean, i.e.
$\sigma=\frac{1}{N-1}\sum_{i=0}^{N-2}\rho_{i+1/2}.$
We define the equidistribution loss function $L_{E}$ by
$L_{E}(\bm{\Theta})=\sum_{i=0}^{N-2}(\rho_{i+1/2}-\sigma)^{2}.$ (21)
Remark 1
In practice, to ensure regularity when $u^{\prime\prime}$ is small we replace the definition of $\rho_{i+1/2}$ in (19) by the regularised version
@@
A value of $\epsilon^{2}=0.1$ works well in practice, and we will use it for all further calculations.
4.2.2 Optimal knots for the interpolating FKS
The optimal knots $k_{i},~{}i=0,\ldots,N-1$, of $L_{I}$ for $\Pi_{1}u$ can be determined semi-analytically by using quadrature applied to (19) in the one-dimensional setting. Similar to Moving Mesh PDEs [HuangRussel], we consider the physical interval $x\in[0,1]$ to be a map from a computational interval $\xi\in[0,1]$ so that $x=x(\xi)$ with $x(0)=0$ and $x(1)=1$. For $i=0,\ldots,N-1$, the $ith$ knot point $k_{i}$ is given by $k_{i}=x(i/(N-1))$. Provided that $N$ is large we can then use the approximations
@@ 4.2.2 Optimal knots for the interpolating FKS
The optimal knots $k_{i}=x(i/(N-1))$ of a piecewise linear interpolating FKS on an equidistributed mesh can then be determined from the solution of (23) using (24). We substitute (22) into (18) to obtain
$L_{I}(\bm{\Theta})\approx\sum_{i=0}^{N-2}\left(\frac{dx}{d\xi}(\xi_{i+1/2})\right)^{5}\frac{(u^{\prime\prime}(x(\xi_{i+1/2})))^{2}}{120(N-1)^{5}}=\sum_{i=0}^{N-2}\frac{D^{5}}{120(N-1)^{5}}=\frac{D^{5}}{120(N-1)^{4}}$

### v2

@@ 4 Training the ReLU NN and the FKS
Having seen the expressivity of the ReLU NN and of the FKS and the formal equivalence of the (shallow) ReLU NN and the FKS approximation in Section 2, in this and the following two sections we consider how to train these to find the best approximation of the given function $u$. In all cases the training will be done through a (first order) gradient-based approach using the Adam optimiser with a suitable initial start and with an appropriate loss function. This procedure is chosen as it is a very common way of training a machine learning network. We find in general that a direct training of all of the parameters is problematic, and that a more effective approach is to divide the training procedure into first training the knots/breakpoints and then the weights of the network. Different loss functions are used in these two cases, and preconditioning is needed for the ReLU NN network. To unify notation, we denote the ReLU NN or the FKS approximation by $y(x)\colon[0,1]\to\mathbb{R}$ and the associated parameter values by $\mathbf{\theta}$. Further, we denote the target function by $u\colon[0,1]\to\mathbb{R}$. We initially consider the loss function $\mathcal{L}_{2}^{2}$ based directly on the mean-squared approximation error of the approximation. This is the default loss function which is typically used in training a NN approximation.
As both $y(x)$ and $u(x)$ are given for all $x\in[0,1]$ we have that the $\mathcal{L}^{2}_{2}$-error of the approximation is given by
@@
In practice, we cannot evaluate (13) for a general function and we have to consider various approximations to it. For this, we consider a set of quadrature points $\{x_{k}\}_{k=0}^{s-1}\subset[0,1]$ for some large parameter $\0s\in\mathbb{N}\backslash\{0\}$ which can either be randomly generated with uniform distribution or regularly distributed over the domain $[0,1]$. Assume that points $\{x_{k}\}_{k=0}^{s-1}\subset[0,1]$ are ordered so that $x_{i}<x_{i+1}$. An approximation to $\mathcal{L}_{2}^{2}$ in (13) is given by
$L_{2}^{2}(\mathbf{\theta})=\sum_{i=0}^{s-1}(x_{i+1}-x_{i})\Big{(}y(x_{i},\mathbf{\theta})-u(x_{i})\Big{)}^{2}.$ (14)
@@
Note that minimising (14) is a highly non-convex problem, and the minimisation process generally leads to sub-optimal solutions. A similar loss function is often used in the Deep Ritz Method (DRM) [8]. Similarly to the DRM, we will consider both the case where $\{x_{k}\}_{k=0}^{s-1}$ are sampled at each iteration or fixed during the training. Observe that the minimisation of $L_{2}^{2}$ in (14) reduces to the pointwise error at sample points $\{x_{i}\}_{i=0}^{s-1}$ and neglects global properties of the function, such as its curvature. This is important when considering equidistributed knots in regions of highest curvature, which contribute significantly to the approximation error.
Whilst simple and widely used, for example in general ReLU NN training or in linear spline approximation, the use of the loss function (14) is problematic when used to train either a ReLU NN or an FKS network. We can see this clearly in the results presented in Figure 1 where we see poor accuracy of both the FKS and ReLU NN networks trained using this loss function. Similar behaviour can be seen in many other examples and we will give a series of numerical calculations giving evidence for this is Section 8.
There are a number of reasons for this poor behaviour. Firstly all of the terms in the parameter set $\mathbf{\theta}$ are treated equally, whereas the knots/breakpoints $k_{i}$ play a very different role in the approximation from the weights $w_{i},c_{i}$. This leads to a highly non-convex minimisation problem. Secondly, in the the case of the ReLU NN approximation it leads (as we shall show) to an ill-conditioned problem for the weights.
Accordingly we now consider an alternative procedure for finding the best approximation which takes into account the distinguished role played by the knots/breakpoints. This firstly solves the nonlinear problem of finding a close approximation to the knots and then the nearly linear problem of finding the wights and the optimal knots of the network. To do this we make use of the interpolating FKS
5 Finding the optimal knots of the interpolating FKS using equidistributon based loss functions
@@
Whilst the interpolating FKS (IFKS) $\Pi_{1}u$ is sub-optimal as a general approximating function, it is an excellent first guess for general optimisation procedures. Finding the IFKS only involves determining the knots $k_{i}$ for $i=1,\ldots,N-1$ which leads to a simplified approximation process approach for choosing the optimal knots. We show presently that this process can be adapted for training the breakpoints of the ReLU  NN. Having found the knots, the problem of finding the weights of the optimal FKS is a well conditioned nearly linear problem and we consider this in the next section together with the problem of finding the coefficients of the ReLU  NN.
To find the optimal knots for the IFKS we consider optimising a different loss function from the standard $L_{2}$ loss above. This loss function is based on the concept of equidistributing the IFKS solution error across the knot locations.
5.1 The equidistribution loss function
@@
$y(x,\mathbf{\theta})-u(x)\approx\frac{1}{2}(x-k_{i})(k_{i+1}-x)u^{\prime\prime}(x_{i+1/2})$ (15)
for any $x\in[k_{i},k_{i+1}]$ and some $x_{i+1/2}\in(k_{i},k_{i+1})$. Note that this result even holds for the functions such as $u(x)=x^{2/3}$ for which $u^{\prime\prime}(0)$ is singular. This yields
@@
$L_{I}(\mathbf{\theta})=\frac{1}{120}\sum_{i=0}^{N-2}\;(k_{i+1}-k_{i})^{5}\;|u^{\prime\prime}(k_{i+1/2})|^{2}$ (17)
where $k_{i+1/2}=k_{i}+k_{i+1}$ and we obtain $\mathcal{L}_{2}^{2}(\mathbf{\theta})\approx L_{I}(\mathbf{\theta})$. Note that $L_{I}$ only depends on the free knot locations for the interpolating FKS, and thus can be used to train the knots. However, unlike other loss functions $L_{I}$ requires knowledge of $u^{\prime\prime}$ rather than point values of $u$ and we will assume for the moment that $u^{\prime\prime}$ is known.
A powerful result in the form of the equidistribution principle, first introduced by de Boor [2] for solving boundary value problems for ordinary differential equations, gives a way of finding $\mathbf{\theta}$ so that $L_{I}$ in (17) is minimised. More precisely, minimising $L_{I}$ in (17) can be expressed by equidistributing the error over each cell which results in the following lemma (see [15, Chapter 2]):
Lemma 5.1 [Equidistribution] The loss $L_{I}$ in (17) is minimised over all knots $k_{i}$ if
@@
The equidistribution principle in Lemma 5.1 replaces a non-convex optimisation problem for finding all of the terms of ${\mathbf{\theta}}$ by one of just finding the knots ${\mathbf{k}}$ with better convexity properties. A set of knot points satisfying these conditions is said to be equidistributed. The degree of equidistribution can then be used as a quality measure of the solution.
The algebraic equations (19) can in principle be solved directly [3], or iteratively using moving mesh techniques such as [14] for example. We propose instead training the knots of the IFKS directly using an optimisation approach enforcing the equidistribution condition (19) directly through minimising an equidistribution-based loss function. We do this using Adam with an initially uniform distribution of the knots as a start. We demonstrate that this procedure is both effective, and can be generalised to the training of the ReLU NN.
To do this, we compute $\rho_{i+1/2}$ from knots $k_{i}$ by (18) and set $\sigma$ as their mean, i.e. $\sigma=\frac{1}{N-1}\sum_{i=0}^{N-2}\rho_{i+1/2}.$ We define the equidistribution loss function $L_{E}$ by
$L_{E}(\mathbf{\theta})=\sum_{i=0}^{N-2}(\rho_{i+1/2}-\sigma)^{2}.$ (20)
Remark 5.2 In practice, to ensure regularity when $u^{\prime\prime}$ is small we replace the definition of $\rho_{i+1/2}$ in (18) by the regularised version
@@
A value of $\epsilon^{2}=0.1$ works well in practice, and we will use it for further calculations.
5.2 Convergence of the IFKS
To show why it is important to first find the knots of the IFKS we show the excellent accuracy ${\mathcal{O}}(1/N^{4})$ of this approximation. The optimal knots $k_{i},~{}i=0,\ldots,N-1$, of $L_{I}$ for $\Pi_{1}u$ are given analytically by using quadrature applied to (18) in the one-dimensional setting, and we will consider this approach in Section 8.3. Similar to Moving Mesh PDEs [15], we consider the physical interval $x\in[0,1]$ to be a map from a computational interval $\xi\in[0,1]$ so that $x=x(\xi)$ with $x(0)=0$ and $x(1)=1$. For $i=0,\ldots,N-1$, the $ith$ knot point $k_{i}$ is given by $k_{i}=x(i/(N-1))$. Provided that $N$ is large we can then use the approximations
@@ 5.2 Convergence of the IFKS
The optimal knots $k_{i}=x(i/(N-1))$ of a piecewise linear IFKS on an equidistributed mesh can then be determined from the solution of (22) using (23). We substitute (21) into (17) to obtain
$L_{I}(\mathbf{\theta})\approx\sum_{i=0}^{N-2}\left(\frac{dx}{d\xi}(\xi_{i+1/2})\right)^{5}\frac{(u^{\prime\prime}(x(\xi_{i+1/2})))^{2}}{120(N-1)^{5}}=\sum_{i=0}^{N-2}\frac{D^{5}}{120(N-1)^{5}}=\frac{D^{5}}{120(N-1)^{4}}$